# Generating Clinically Realistic EHR Data via a Hierarchy- and Semantics-Guided Transformer

### v1

@@ 2.2 Utility Evaluation
To evaluate the clinical utility of synthetic EHR data, we conduct a disease classification experiment that assesses whether predictive models trained on synthetic clinical data can generalize to real-world patient records. This experiment follows the Train on Synthetic, Test on Real (TSTR) evaluation paradigm, where models are first trained on synthetic data and then evaluated on real patient outcomes. This methodology ensures that synthetic data captures meaningful clinical patterns that can support downstream predictive modeling tasks (Esteban et al., 2017). In our setting, we adopt the phenotype classification task, considering 25 phenotypes, a widely used benchmark in clinical predictive modeling (Harutyunyan et al., 2019). For real data, phenotype labels are assigned using a rule-based mapping from ICD diagnosis codes to phenotype categories, based on the HCUP Clinical Classifications Software (CCS) (Harutyunyan et al., 2019). Each patient in the real dataset is assigned one or more phenotype labels based on their recorded ICD codes. For synthetic data, phenotype labels are first generated by the model and then used to conditionally generate patient visit sequences containing medical codes, following the approach in (Esteban et al., 2017). This conditional generation strategy ensures that patient trajectories are synthesized in a way that maintains meaningful associations between phenotype labels and clinical events. For the classification model, we employ a bi-directional LSTM classifier with a fully connected output layer (Theodorou et al., 2023). For each dataset, we randomly extract 5,000 patient records for training, maintaining a balanced distribution of positive and negative labels for every single chronic disease category. Further, we reserve a separate validation set of 250 records for model selection and 500 real patient records as a held-out test set. Table 2 presents the classification results for various baseline models. We report accuracy, precision, recall, and F1-score, with standard deviations computed over the 25 phenotype labels. The performance of models trained on real data serves as an upper bound for comparison. Among synthetic data models, HiSGT consistently achieves the highest performance, surpassing HALO and other baseline methods in both MIMIC-III and MIMIC-IV datasets. These results highlight the effectiveness of our conditional generation strategy, demonstrating that HiSGT-generated patient trajectories retain the clinical coherence necessary for accurate phenotype classification. Table 3 provides a detailed breakdown of classification performance across different disease types (acute, chronic, and mixed). The macro-averaged accuracy and F1-score across each disease category offer insights into how well synthetic data generalizes across different clinical conditions. Across all disease types, HiSGT demonstrates superior performance over current SOTA method HALO.
2.3 Privacy Evaluation
In addition to ensuring high fidelity and utility, it is critical that synthetic EHR data do not compromise patient privacy (Giuffrè and Shung, 2023). To this end, we evaluate the privacy protection of our synthetic data by simulating two common attacks: a membership inference attack (MIA) and an attribute inference attack (AIA). These attacks are designed to assess whether an adversary could either determine if a particular patient record was used during training or infer sensitive attributes from the synthetic records. Specifically, MIA tests whether an adversary can correctly classify a real patient record as having been used (or not used) during the synthetic data generation process. We construct an attack dataset by randomly sampling 500 records from the training set (labeled as positive) and 500 records from the test set (labeled as negative). For each record in this attack set, we compute the minimum Hamming distance to the synthetic dataset. By comparing each distance to the median distance of the entire attack set, records are classified as members (positive) if the distance is below the median and as non-members otherwise. An accuracy close to 50% indicates that the synthetic data do not leak membership information beyond what would be expected by random guessing. The AIA aims to determine whether sensitive attributes can be reliably predicted from the synthetic data. We transform both real and synthetic patient records by aggregating visit-level information and then select a set of common codes (derived from the most frequent 100 codes in the training data) to serve as a basis for inference. For each record, we compare the known label set with those of its nearest neighbors (based on a simple set-based distance) and predict additional codes via a majority vote. The performance of the attack is measured using an F1-score (AIA-F1), where lower scores indicate better privacy protection by the synthetic data. Table 4 summarizes the results of both MIA and AIA on MIMIC-III and MIMIC-IV. The reported metrics include accuracy, precision, recall, and F1-score for MIA, as well as the AIA-F1 score. As evidenced by the results, all evaluated models—including HiSGT—exhibit an attack accuracy near 50% and low AIA-F1 scores, which confirms that the synthetic data do not reveal substantial information about individual patient membership or sensitive attributes. In particular, Table 4 reveals an inherent trade-off between fidelity and utility with privacy in synthetic EHR generation. Models such as LSTM, EVA, and SynTEG achieve the most robust privacy protection. Their MIA accuracies are around 50%—consistent with random guessing—and they report the lowest AIA-F1 scores (e.g., 0.007–0.008 on MIMIC-III), indicating that these methods leak minimal information regarding patient membership and sensitive attributes. However, this strong privacy comes at a cost: these same methods exhibit considerably lower fidelity and utility (as shown in Tables 1 and 2), meaning that their synthetic data deviate substantially from the true data distributions and present limited utility. Conversely, SOTA methods like HALO and the proposed HiSGT deliver higher fidelity and utility. When compared to HALO, HiSGT achieves similar privacy protection with an MIA F1-score of 0.472 and an AIA-F1 of 0.034 on MIMIC-III, while continuing to improve fidelity and utility.
Table 4: Privacy Evaluation Results. This table reports the performance of the membership inference attack (MIA) and attribute inference attack (AIA) on both MIMIC-III and MIMIC-IV. The metrics include accuracy, precision, recall, and F1-score for MIA, and the AIA-F1 score for AIA. An accuracy near 50% indicates that the synthetic data do not leak membership information beyond chance levels, while low AIA-F1 scores reflect robust protection against inferring sensitive attributes. Notably, the results highlight a trade-off between fidelity and utility with privacy: while models such as LSTM, EVA, and SynTEG achieve very strong privacy (evidenced by the lowest AIA-F1 scores), they suffer from poor fidelity and utility. In contrast, SOTA methods like HALO—and our proposed HiSGT—strike a better balance by delivering higher fidelity and utility with only a moderate increase in privacy risk.
@@ 2.3 Privacy Evaluation
2.4 Interpretable Input Embeddings via t-SNE Visualizations
To further evaluate the ability of our model to capture meaningful hierarchical and semantic information, we perform t-SNE visualizations of input embeddings after training HiSGT and use HALO as a comparison. These visualizations help assess whether the model successfully groups medical codes into coherent clusters. From Figure 3, we observe that HALO exhibits more scattered clusters, with many points overlapping across different ICD categories. This suggests that HALO’s embedding space does not strongly enforce category-specific separation. In contrast, HiSGT demonstrates more defined and compact clusters, indicating that it captures category-level information more effectively. The improved clustering structure also leads to a more interpretable embedding space. These findings indicate that the incorporation of hierarchical and semantic representations in HiSGT enhances the clinical coherence of medical code embeddings. By preserving these relationships more effectively, HiSGT has the potential to improve fidelity and utility in synthetic EHR generation, since the generated sequences better reflect real-world clinical distributions.
(a) HALO
(b) HiSGT
Figure 3: t-SNE visualizations of input embeddings. Each point represents a medical code, colored by ICD category. HALO (left) and HiSGT (right) show distinct clustering behaviors, where HALO presents more scattered clusters while HiSGT exhibits more compact and category-level clusters. The legend represents the first character of ICD chapters, e.g., ‘R’ corresponds to ‘Symptoms, signs and abnormal clinical and laboratory findings, not elsewhere classified’.
3 Discussion
In this study, we presented HiSGT, a novel Hierarchy- and Semantics-Guided Transformer for generating synthetic EHR data with high clinical fidelity. By integrating hierarchical information from clinical coding taxonomies and semantic representations from pre-trained clinical language models, HiSGT overcomes key limitations of prior methods that treat medical codes as isolated tokens. Our extensive evaluations on the MIMIC-III and MIMIC-IV datasets demonstrate that HiSGT produces synthetic patient records that more accurately replicate real-world clinical distributions while maintaining robust privacy and competitive utility. The core fidelity metrics reveal that HiSGT substantially improves the preservation of both marginal event distributions and intra- and inter-visit dependencies compared to baseline methods. For instance, our model achieved an 8.2% improvement in the intra-visit Bigram score over HALO, indicating better capturing of the co-occurrence patterns among clinical events. These improvements are particularly important because accurate modeling of such dependencies is crucial for ensuring that synthetic data support downstream clinical applications. Further, HiSGT is a fully Transformer-based model and, contrary to HALO, does not rely on a dense layer to capture intra-visit correlations (with parameters that scale quadratically with vocabulary size), making it easier to scale to very large code vocabularies. Utility evaluations further confirm that HiSGT-generated data enable the training of disease classification models that generalize well to real patient records. Although our experiments indicate that the classification performance using synthetic data is on par with SOTA baselines like HALO, HiSGT achieves this while also exhibiting higher fidelity. In our setting, the synthetic labels are generated as part of the conditional process, ensuring that the joint distribution of features and labels is preserved. This self-contained generation process minimizes the need for post-hoc label assignment and contributes to a more natural alignment between the synthetic records and their clinical interpretations. The privacy evaluation underscores a crucial trade-off between fidelity, utility, and privacy. Models such as LSTM, EVA, and SynTEG offer strong privacy protection—as evidenced by near-chance membership inference accuracy and low attribute inference F1-scores—but suffer from poor fidelity and limited utility. In contrast, HiSGT strikes an optimal balance by higher fidelity and utility with only a moderate increase in privacy risk. This balance is particularly valuable for clinical applications where preserving the nuanced relationships in patient data is as important as ensuring confidentiality. While this study focuses on integrating hierarchical and semantic representations for ICD diagnosis codes, the proposed paradigm is inherently generalizable to other medical code type including procedure and medication codes.
To evaluate the contributions of hierarchical embeddings, semantic embeddings, and consistency constraints, we conduct an ablation study (Table 5) to help analyze the different components in HiSGT. Starting with a base Transformer trained solely on medical codes, without structured knowledge, we observe that it struggles to capture co-occurrence and longitudinal dependencies. Introducing hierarchical or semantic embeddings leads to gains in all four metrics, especially Bigram and Seq Bigram. The combined use of hierarchical and semantic embeddings yields further improvements. We notice a concurrent approach, MEDTOK (Su et al., 2025), that also explores relational structure and text descriptions for medical code representations. However, its methodology differs significantly: MEDTOK trains a tokenizer with multiple loss functions to improve code embeddings for downstream tasks, whereas HiSGT not only encodes hierarchical and semantic knowledge but also enforces consistency constraints during generation. In our ablation study, the closest equivalent to MEDTOK is the ‘+ Hierarchical & Semantic Embds’ variant. While this model improves fidelity compared to a base Transformer, it remains inferior to HiSGT’s full model due to the lack of consistency regularization.
Table 5: Ablation Study of HiSGT Components. Each row adds a different form of embedding or consistency constraint to a base Transformer. A concurrent work, MEDTOK (Su et al., 2025), (like our variant ‘+ Hierarchy & Semantic Embds’) improves upon the code-only baseline but cannot match our full model’s fidelity, highlighting the importance of consistency regularization.
@@ 3 Discussion
+ Both Embds & Consistencies 0.984 0.949 0.879 0.976 0.989 0.967 0.948 0.989
Our Full HiSGT
Despite these promising results, several limitations warrant further investigation. First, while HiSGT successfully integrates hierarchical and semantic information, its performance is still partially constrained by the quality and granularity of the underlying clinical ontologies and pre-trained language models. For example, even though we employ ClinicalBERT—which is pre-trained on the MIMIC dataset—it is still limited in capturing the precise semantic context. Our preliminary attempts to use contrastive learning to refine the semantic embeddings incurred significant computational overhead while yielding only marginal improvements. Future work should consider exploring alternative strategies or more domain-specific pre-training to further enhance semantic representation. Second, our utility evaluation was based on a multi-class classification paradigm for disease classification. While this task is representative, it may not fully capture the diverse clinical decision-making processes. Evaluating HiSGT on a broader range of downstream tasks—such as risk stratification or temporal event prediction—could provide a more comprehensive assessment of its clinical utility. Finally, scalability poses an important challenge. As the complexity of EHR data increases—with richer feature sets and longer patient histories—the computational cost of incorporating detailed hierarchical and semantic representations may become prohibitive. Even after optimizing the model architecture to include the most essential embeddings and consistency regularization terms, scalability remains a potential issue. Future research should focus on optimizing training strategies and exploring more efficient model architectures to ensure that HiSGT can be applied effectively in large-scale settings.
In conclusion, our findings suggest that HiSGT represents a significant step forward in synthetic EHR data generation by balancing high clinical fidelity, robust privacy protection, and practical utility. By leveraging structured domain knowledge through hierarchical and semantic embeddings, HiSGT not only produces synthetic data that closely mirror real patient records but also facilitates downstream clinical modeling tasks. Importantly, the use of hierarchically and semantically meaningful code embeddings will also lead to more meaningful patient representations when codes are aggregated, e.g. for clustering purposes. This work opens new avenues for developing advanced synthetic data generators that can ultimately support a wide range of healthcare applications—from model benchmarking and data augmentation to collaborative research in settings where access to real patient data is limited.
4 Methods
4.1 Problem Formulation
Electronic health record data naturally exhibits a multi-level structure comprising patients, visits and clinical events. Let $\mathcal{X}$, $\mathcal{V}$ and $\mathcal{E}$ denote the sets of patients, visits and clinical events, respectively. For each patient $\bm{x}\in\mathcal{X}$, the medical history is represented as a sequence of visits $\bm{x}=\{\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{n_{x}}\}$, where $n_{x}$ is the number of visits. Each visit $\bm{v}_{i}$ comprises a set of medical codes representing diagnoses, procedures, and other clinical events: $\bm{v}_{i}=\{e^{i}_{1},e^{i}_{2},\ldots,e^{i}_{m_{i}}\},$ with $m_{i}$ denoting the number of medical codes recorded during visit $\bm{v}_{i}$. The objective of synthetic EHR generation is to model the joint distribution of patient trajectories while preserving both intra-visit (within-visit) and inter-visit (longitudinal) dependencies. The model estimates the conditional probability of each clinical event given past visits and previously observed events within the same visit:
where $\Theta$ represents the model parameters. The generative model seeks to maximize this likelihood to ensure that the generated sequences closely resemble real-world clinical patterns.
Once trained, a generative model can synthesize realistic EHR trajectories by predicting medical events based on past events:
$e_{j}^{i}\sim P\Bigl{(}e^{i}_{j}\,\Big{|}\,\bm{v}_{1},\cdots,\bm{v}_{i-1},e^{i}_{1},\cdots,e^{i}_{j-1}\Bigr{)}$ (2)
At each step $j$, the predicted event $e_{j}^{i}$ is appended to the sequence, and the updated sequence is fed back into the model to generate the next event. The process continues until a predefined maximum length $L$ is reached or an ‘END_RECORD’ token is generated, signaling the completion of a synthetic patient record.
4.2 Dataset and Baseline Details
We conduct experiments on two widely used real-world EHR datasets: MIMIC-III (Johnson et al., 2016) and MIMIC-IV (Johnson et al., 2023). To construct patient-level longitudinal records, we first extract hospital admissions and ICD-9 diagnosis codes from both datasets. Admissions are sorted chronologically based on timestamps to maintain the correct temporal sequence of patient visits. Diagnosis codes are then aggregated at the hospital visit level, and visits are grouped by patient index, forming structured patient records that capture the progression of medical events over time. For utility evaluation, we map ICD codes to phenotype categories using a predefined benchmark taxnomy (Harutyunyan et al., 2019). Each patient is assigned one or more phenotype labels (up to 25 labels) based on their recorded diagnoses. The dataset is then tokenized by constructing a vocabulary that includes all unique ICD codes in the dataset, 25 phenotype label tokens, and five special tokens used for sequence structuring. These special tokens: START_RECORD indicates the beginning of a patient record, END_LABEL denotes the transition from phenotype labels to visit sequences, END_VISIT marks the end of a hospital visit, END_RECORD signifies the completion of a patient record, and ‘PADDING’ ensures standardized sequence lengths. After preprocessing, MIMIC-III consists of 46,520 patients with 6,984 unique ICD diagnosis codes and 7,012 unique tokens, while MIMIC-IV includes 124,525 patients with 9,072 unique ICD codes and 9,102 unique tokens. Both datasets contain structured longitudinal patient records with multiple hospital visits, making them well-suited for evaluating synthetic EHR generation models.
To assess the effectiveness of HiSGT, we compare it against six state-of-the-art (SOTA) generative models used for EHR synthesis. LSTM (Lee, 2018) is a recurrent neural network that models sequential dependencies. GPT (Radford et al., 2019) applies a transformer-based autoregressive approach for next-token prediction. EVA (Biswal et al., 2021) uses a variational autoencoder (VAE) to learn latent representations of EHR sequences. SynTEG (Zhang et al., 2021) employs a GAN-based framework to generate synthetic patient visits. HALO-Coarse (Theodorou et al., 2023) treats each visit as a single token, modeling longitudinal dependencies at a coarse level, while HALO (Theodorou et al., 2023) refines this approach by representing visits at a more granular level.
4.3 Experimental Setting
We conduct experiments using the MIMIC-III v1.4 and MIMIC-IV v2.2 datasets. The data is split into 80%-20% for training and testing, with an additional 90%-10% split within the training set for validation. For HiSGT, we use a six-layer Transformer architecture with eight attention heads and a hidden size of 384. Training is performed using the Adam optimizer with a learning rate of $10^{-4}$, a batch size of 48, and a dropout rate of 0.1. The input sequence length is set to 768 tokens, and training is run for 100 epochs with early stopping (patience = 10) based on validation loss. The best model checkpoint is saved according to minimum validation loss. The model is implemented in Python 3.13.0 using PyTorch 2.5.0+cu121, along with scikit-learn 1.5.2, NumPy 2.1.2, and transformers 4.46.2. All experiments are conducted on a single NVIDIA H100 GPU (80GB VRAM). The HiSGT source code is publicly available at: https://github.com/jameszhou-gl/HiSGT.
4.4 HiSGT Framework
4.4.1 Tokenization
The tokenization process in HiSGT structures raw EHR data into a sequence of discrete tokens while preserving temporal dependencies betweeen medical events. We extract longitudinal patient records from the MIMIC-III and MIMIC-IV datasets, using key tables such as patients.csv, admissions.csv and diagnoses_icd.csv. Each patient’s health timeline consists of one or more hospital visits, with each visit containing diagnosis codes that describe the patient’s conditions. To ensure an accurate representation of patient trajectories, the raw data undergoes preprocessing, where hospital visits are first sorted chronologically to maintain the correct sequence of admissions. Diagnosis codes are then grouped by patient ID, ensuring that all events for a single patient remain together. Additionally, each patient is assigned phenotype labels derived from their diagnoses, according to predefined clinical mappings. Once preprocessed, the dataset is tokenized by constructing a vocabulary that includes all unique ICD codes, along with phenotype label tokens and special tokens indicating sequence boundaries. As shown in Figure 1(a), each patient sequence follows a structured format. It begins with a START_RECORD token, marking the initialization of a new patient record. This is followed by phenotype label tokens that summarize the patient’s conditions, such as Hypertension with complications, Pneumonia and Chronic kidney disease. These label tokens play a crucial role in conditioning the generation of subsequent visits, ensuring consistency with standard utility evaluations. The END_LABEL token separates phenotype labels from individual hospital visits. Each visit is represented as a collection of medical event tokens corresponding to ICD codes, e.g., ICD9-40301 (Hypertensive chronic kidney disease) and ICD9-486 (Pneumonia, organism unspecified). To preserve temporal dependencies, events from the same visit are grouped and concluded with an END_VISIT token. The sequence may progress through multiple visits, maintaining the correct ordering of hospital admissions. Finally, the END_RECORD token indicates the end of a patient’s medical history. Since patient trajectories vary in length, PADDING tokens are appended where necessary to ensure consistent input dimensions.
4.4.2 Hierarchical & Semantic Embeddings
Semantic Embeddings. Beyond hierarchical relationships, medical codes are accompanied by textual descriptions that provide additional clinical context. These descriptions are essential for distinguishing between related conditions, procedures, and medications. For example, the ICD-10 code ‘E11.65’ specifies ‘Type 2 diabetes mellitus with hyperglycemia’ whereas ‘Z13.6’ specifies ‘screening for cardiovasscular disorders’. Ignoring these textual descriptions can lead to synthetically generated records that fail to reflect meaningful clinical distinctions. HiSGT employs ClinicalBERT (Alsentzer et al., 2019), a pre-trained clinical language model, to encode these descriptions into dense semantic representations. Given the text description $T(e)$ corresponding to medical code $e$, ClinicalBERT generates a semantic embedding: $\bm{s}_{e}=\mathrm{BERT}(T(e))$. Since ClinicalBERT outputs embeddings in a fix and high-dimensional space, we apply a linear transformation to project them into the model’s latent space: $\bm{Z}_{s}=\bm{W}_{s}\bm{s}_{e}+\bm{b}_{s}$; where $\bm{W}_{s}\in\mathbb{R}^{d_{s}\times d}$ and $\bm{b}_{s}\in\mathbb{R}^{d}$ are learnable parameters.
$\bm{Z}\;=\;\bm{Z}_{t}\;+\;\bm{Z}_{h}\;+\;\bm{Z}_{s}$ (3)
4.4.3 Model Structure
HiSGT follows a decoder-only Transformer architecture (Radford et al., 2019), generating synthetic EHR sequences in an auto-regressive (AR) manner. The model processes tokenized sequences and predicts the next medical event based on prior context. As shown in Figure 1(c), the input to HiSGT consists of three components: a code token embedding $\bm{Z}_{t}$, a hierarchical embedding $\bm{Z}_{h}$, and a semantic embedding $\bm{Z}_{s}$. These embeddings are summed together to form the input embedding $\bm{Z}$ and positional embeddings are added further to ensure the model understands the sequential nature of patient trajectories. The processed embeddings pass through six stacked Transformer decoder blocks, each comprising a masked multi-head self-attention mechanism and a feed-forward network. The attention mechanism selectively attends to relevant past events while maintaining temporal dependencies across hospital visits. A causal mask ensures that the model generates future events based only on past visits, preserving a realistic medical progression. Following the self-attention layer, the feed-forward network refines token representations through nonlinear transformations. The output of each decoder block is iteratively refined through multiple layers to progressively learn contextual dependencies. At the final stage, the hidden states of the Transformer blocks are projected onto different output heads. The next medical event in the sequence is predicted through the code head, which computes probabilities over the medical code vocabulary. To ensure that generated sequences remain clinically structured and meaningful, HiSGT introduces two additional regularization heads. The hierarchical consistency head predicts a hierarchy-aware representation, reinforcing the structured taxonomic relationships within the ICD system. The semantic consistency head maps generated medical codes back into the semantic space, ensuring that their meaning aligns with real-world clinical descriptions. While the primary objective is next-code prediction, the hierarchical and semantic consistency heads serve as essential constraints that guide the model toward more realistic and interpretable patient trajectories.
4.4.4 Training and Inference
HiSGT is trained using a multi-objective optimization approach that balances next-code prediction with hierarchical and semantic consistency constraints. The next-code prediction task is framed as a standard cross-entropy loss, where the model learns to predict the next code given the prior sequence of tokens. Besides, we introduce two auxiliary loss terms to ensure that generated sequences adhere to real-world medical knowledge. HiSGT employs two lightweight auxiliary heads: a hierarchy head and a semantic head. The hierarchy head maps the hidden state of each generated token back to the hierarchical embedding space, ensuring that medical codes respect their taxonomic relationships. Similarly, the semantic head projects hidden states onto the semantic embedding space, aligning generated codes with clinically meaningful representations obtained from ClinicalBERT. Both consistency objectives are formulated as mean squared error (MSE) losses, minimizing the discrepancy between predicted embeddings and their original input counterparts. These auxiliary heads act as implicit regularization mechanisms, guiding the model toward generating sequences that maintain both structural coherence and semantic fidelity. Once trained, HiSGT generates synthetic patient trajectories in an AR manner. Given an initial START_RECORD token, the model first generates phenotype labels, summarizing the patient’s overall conditions. Upon detecting the special token END_LABEL, it proceeds to generate medical codes for each visit while preserving temporal dependencies. The process continues until the model outputs the END_RECORD token, marking the valid completion of a synthetic patient sequence.
4.5 Comparison of HiSGT to Related Work
EHR Data Generation. The generation of synthetic EHRs have evolved across several categories (Chen et al., 2024). Eearly rule-based methods rely on predefined clinical guidelines to simulate patient records (Buczak et al., 2010; Franklin et al., 2014; McLachlan et al., 2018). While these methods ensure clinically interpretability, they fail to capture complex correlations in real-world EHR data. GAN-based approaches model the distribution of real EHR data through adversarial training. A generator synthesizes patient records, while a discriminator distinguishes between real and synthetic samples (Zhang et al., 2020; Yale et al., 2020; Wang et al., 2024; Li et al., 2023; Baowaly et al., 2019; Yang et al., 2019; Lee et al., 2020; Yan et al., 2021; Torfi and Fox, 2020; Kuo et al., 2022). VAE-based methods learn a latent representation of EHR data and generate new samples by sampling from the learned distribution (Biswal et al., 2021; Sun et al., 2024). However, both of GANs and VAEs struggle with rare events in high-dimensional EHR data and capturing fine-grained longitudinal dependencies in EHR data. Recently, Transformer-based generative models gained traction due to their ability to capture long-range dependencies in patient records. HALO (Theodorou et al., 2023) applies a self-attention mechanisms to model the sequential structure of medical codes, significantly expanding the vocabulary of synthetic data from fewer than 100 tokens to over thousands. However, existing methods rely solely on modeling raw code sequences. As result, the generated records lack high clinical fidelity. In this work, we propose a Transformer-based framework with two inductive biases of hierarchical and semantic information.
Generative Transformer. The Transformer architecture, introduced by  (Vaswani et al., 2017), has revolutionized sequence modeling by effectively capturing long-range dependencies through self-attention mechanisms. In natural language processing (NLP), generative Transformer models such as GPTs series (Radford et al., 2019; Brown et al., 2020) have demonstrated SOTA performance in text generation. These models predict tokens sequentially while conditioning on all previous tokens, making them highly effective for learning complex dependencies. Beyond NLP, Transformers have been successfully adapted to other domains including computer vision (Dosovitskiy et al., 2021) and multimodal learning (Hurst et al., 2024; Yang et al., 2023). In healthcare, Transformer-based models for synthetic EHR generation typically treat patient trajectories as sequences of discrete medical codes (Theodorou et al., 2023; Pang et al., 2024). Although they enable scaling to synthetic data with thousands of unique tokens, they model medical codes purely as discrete symbols, thereby the generated EHR lack clinical fidelity. Our work extends Transformer-based EHR generation by explicitly incorporating structured knowledge into the generative process—leveraging hierarchical representations and semantic embeddings inherent in clinical coding systems.
Additional Information
Data Availability
The MIMIC-III (Johnson et al., 2016) and MIMIC-IV (Johnson et al., 2023) electronic health record (EHR) datasets used in this study are publicly available. Access to these datasets requires completion of the necessary training and credentialing process via PhysioNet. Detailed information on obtaining access can be found at https://physionet.org.
Code Availability
The source code for all experiments, including dataset preprocessing, model implementation, training procedures, synthetic data generation, and evaluation, is publicly available at https://github.com/jameszhou-gl/HiSGT.

### v2

@@ 4.3.2 Utility Evaluation
Table 4: Privacy Evaluation Results: performance of the membership Inference Attack (MIA) and Attribute Inference Attack (AIA) on MIMIC-III and MIMIC-IV. Accuracy near 50% and low AIA-F1 scores indicate better privacy.
@@
We evaluate clinical utility using a Train on Synthetic, Test on Real (TSTR) evaluation paradigm, where models are first trained on synthetic data and then evaluated on real patient outcomes [11]. We adopt the 25-phenotype classification benchmark [15]. For real data, labels are assigned using a rule-based mapping from ICD-to-phenotype mapping (HCUP Clinical Classifications Software) [15]. Each patient in the real dataset is assigned one or more phenotype labels based on their recorded ICD codes. For synthetic data, labels are generated first, followed by conditional generation of patient visit sequences, ensuring meaningful label-event associations [11]. We employ a bi-directional LSTM classifier [33]. For each dataset (MIMIC-III/IV), we randomly extract 5,000 patient records for training, maintaining a balanced distribution of positive and negative labels for every single chronic disease category. Further, we reserve a separate validation set of 250 records for model selection and 500 real patient records as a held-out test set.
Table 2 shows TSTR results. Models trained on real data serves as an upper bound. Among synthetic methods, HiSGT consistently achieves the highest performance, surpassing HALO and other baseline methods on both datasets. This demonstrates that HiSGT-generated patient trajectories retain the clinical coherence necessary for accurate phenotype classification. Furthermore, detailed results across disease types in Table 3 confirm HiSGT’s superior generalization compared to SOTA method-HALO for acute, chronic, and mixed conditions.
4.3.3 Privacy Evaluation
To ensure synthetic data do not compromise patient privacy [13], we evaluate protection against two common attacks: Membership Inference Attack (MIA) and Attribute Inference Attack (AIA). The MIA tests if an adversary can identify records used during training. We follow a standard setup: randomly sample 500 records from the training set as members (positive) and 500 from the test set (non-members). For each record, we compute its minimum Hamming distance to the synthetic dataset and classify it as a member if its distance is below the median distance of the attack set. An accuracy near 50% indicates privacy preservation close to random guessing. The AIA assesses if sensitive attributes (medical codes) can be inferred from synthetic data. We aggregate visit-level information and use the 100 most frequent training codes as potential attributes. For each record, we predict undisclosed codes using a majority vote from its nearest neighbors (based on set distance) in the synthetic data. Attack performance is measured by F1-score (AIA-F1), with lower scores indicating better privacy.
Table 4 summarizes the results. All evaluated models, including HiSGT, exhibit MIA accuracy close to 50% and low AIA-F1 scores, confirming reasonable protection against these attacks. In particular, the results highlight an inherent privacy-utility trade-off: models like LSTM, EVA, and SynTEG achieve the strongest privacy scores (lowest AIA-F1, such as 0.007–0.008 on MIMIC-III) but suffer from poor fidelity and utility (as shown in Tables 1 and 2). Conversely, SOTA methods like HALO and our HiSGT deliver higher fidelity and utility. When compared to HALO, HiSGT achieves similar privacy protection with an MIA F1-score of 0.472 and an AIA-F1 of 0.034 on MIMIC-III, while continuing to improve fidelity and utility.
Table 5: Ablation Study. This table presents the impact of integrating hierarchical embeddings, semantic embeddings, and consistency constraints into the base Transformer model.
@@ 4.3.3 Privacy Evaluation
Full HiSGT 0.984 0.949 0.879 0.976 0.989 0.967 0.948 0.989
Figure 2: t-SNE visualizations of input embeddings. Medical codes are plotted based on their input representations for (a) HALO and (b) HiSGT, colored by ICD category. (a) HALO’s standard embeddings show poor separation between categories, forming an undifferentiated cloud. (b) HiSGT’s embeddings, incorporating hierarchical and semantic information, demonstrate substantially improved structure with clearer, more compact, category-specific clusters, reflecting a more clinically coherent representation. The legend represents the first character of ICD chapters [1], e.g., ‘R’ corresponds to ‘Symptoms, signs and abnormal clinical and laboratory findings, not elsewhere classified’.
4.4 Ablation Study
To assess the individual contributions of HiSGT’s components, we conduct an ablation study, systematically evaluating the impact of hierarchical embeddings, semantic embeddings, and consistency constraints on model performance (see Table 5). Starting with a base Transformer trained solely on medical codes (Code Only), we observe reasonable Unigram (0.956) and DimWise (0.964) scores, but it struggles with lower Bigram (0.898) and Seq Bigram (0.821) scores. Introducing either hierarchical embeddings (+ Hierarchical Embd) or semantic embeddings (’+Semantic Embd) alone leads to gains, particularly improving Bigram scores to 0.914 and 0.913, respectively, and Seq Bigram to 0.845 and 0.836. Combining both hierarchical and semantic embeddings (+ Hierarchical & Semantic Embds) yields further improvements across all metrics, notably pushing Bigram to 0.927 and Seq Bigram to 0.859. This demonstrates the benefit of integrating both structural and semantic knowledge into the input representation. Finally, adding the consistency regularization heads on top of both embeddings (Full HiSGT) provides the most significant boost. The full model achieves the best Bigram score (0.949, a substantial increase from 0.927 without consistency) and Seq Bigram score (0.879, up from 0.859). This highlights that both the enriched input embeddings and the consistency constraints guiding the generation process are crucial for HiSGT’s superior fidelity.
4.5 Interpretable Input Embeddings via t-SNE Visualizations
To further evaluate the ability of our model to capture meaningful hierarchical and semantic information, we perform t-SNE visualizations of input embeddings in HiSGT and use HALO as a comparison. These visualizations help assess whether the model successfully groups medical codes into coherent clusters. From Figure 2, we observe that HALO exhibits more scattered clusters, with many points overlapping across different ICD categories. This suggests that HALO’s embedding space does not strongly enforce category-specific separation. In contrast, HiSGT demonstrates more defined and compact clusters, indicating that it captures category-level information more effectively. The improved clustering structure also leads to a more interpretable embedding space. This likely underpins the model’s improved performance in capturing intra- and inter-visit code relationships (Table 1), ultimately contributing to the generation of more realistic synthetic EHR sequences.
5 Conclusion
In this study, we introduced HiSGT, a novel Hierarchy- and Semantics-Guided Transformer for synthetic EHR data generation. By leveraging structured domain knowledge through hierarchical and semantic embeddings, HiSGT generates synthetic data that closely mirror real patient records and facilitates downstream clinical classification tasks. Importantly, the use of hierarchically and semantically meaningful code embeddings will contribute to more interpretable patient representations. While promising, the efficacy of HiSGT is partially constrained by the quality of the external knowledge sources (ontologies and pre-trained LMs). Additionally, assessing HiSGT’s effectiveness across a broader range of downstream application, such as risk stratification or temporal event prediction, and extending HiSGT to handle non-ICD-coded data remain important directions for future work.